# Degenerate parabolic equations in divergence form: fundamental solution and Gaussian bounds

### v2

@@ 6.1. The Cauchy problem and the generalized fundamental solution
Step 1: regularizing coefficients of $A$: We follow [AN24]. Let $\theta\in\mathcal{D}(\mathbb{R})$ a nonnegative function with $\int_{\mathbb{R}}\theta(t)\mathrm{d}t=1$. For all $p\geq 1$, let $\theta_{p}(t)=p\theta(pt)$ be the associated mollifying sequence. We set $A_{p}(t,x):=(\theta_{p}\star A(\cdot,x))(t)$, i.e., we mollify the matrix-valued function A in the time variable only. For all $p\geq 1$ and $t\in\mathbb{R}$, we set
@@
$\partial_{t}\left|U_{p}(t)\right|=\left\{\begin{matrix}\partial_{t}U_{p}(t)\ \mathrm{if}\ U_{p}(t)\geq 0,\\ -\partial_{t}U_{p}(t)\ \mathrm{if}\ U_{p}(t)<0,\\ \end{matrix}\right.\ \ \ \text{and}\ \ \nabla_{x}\left|U_{p}(t)\right|=\left\{\begin{matrix}\nabla_{x}U_{p}(t)\ \mathrm{if}\ U_{p}(t)\geq 0,\\ -\nabla_{x}U_{p}(t)\ \mathrm{if}\ U_{p}(t)<0.\\ \end{matrix}\right.$
@@
Combining Caccioppoli inequality in Lemma 4.3, a weighted Sobolev inequality [HKM18, Theorem 15.26] and the Moser’s iteration principle, we have the following $L^{\infty}$-estimate on nonnegative local weak solutions. For a proof, one can follow the classical scheme or see [Ish99, Proposition 2.1] with lower order coefficients equal to zero.
@@
Moreover, $u\in C([0,\mathfrak{T}];L^{2}_{\omega}(\mathbb{R}^{n}))$ with $u(0)=\psi$, $\lim_{t\to\infty}u(t)=0$ if $\mathfrak{T}=\infty$ (by convention, set $u(\infty)=0$), $t\mapsto\|u(t)\|^{2}_{2,\omega}$ is absolutely continuous on $[0,\mathfrak{T}]$ and we can write the energy equalities. Furthermore, we have $(-\Delta_{\omega})^{\alpha/2}u\in L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))$ for any $\alpha\in(0,1]$ with $r=\frac{2}{\alpha}\in[2,\infty)$ with
@@
Setting $g=0$, the case $\mathfrak{T}<\infty$ reproves [AN24, Theorem 1.2], and we even have a larger uniqueness space.
@@ 6.2. Additional properties using regularity theory for weak solutions
Finally, if $\mathcal{O}=\mathbb{R}^{n}$ and $-\infty<\mathfrak{T}$, then for all $t>s>\mathfrak{T}$ and $x,y\in\mathbb{R}^{n}$, one has
@@
Using (6.4), we have, for all $10\leq i\leq N-$ 1, $u(\tau_{i},z_{i})\leq Cu(\tau_{i+1},z_{i+1})$. Thus, by iterating, we get
@@
for some $\delta>0$ depending only on the structural constants, all $t>s$ and almost every $x,y,h\in\mathbb{R}^{n}$ such that $2\left|h\right|\leq(t-s)^{1/2}+\left|x-y\right|$. For proofs and more details, we refer to [AN24] and [Tru68].
@@
$u(s+\frac{1}{2}(t-s),\tilde{z})$ $=\frac{1}{\omega_{\frac{\gamma}{2}(t-s)}(y)}\int_{B(y,\sqrt{\frac{\gamma}{2}(t-s)})}u(s+\frac{1}{2}(t-s),l)\ \mathrm{d}\omega(l)$
@@
$\tilde{A}(\sigma,z)=\left\{\begin{matrix}A^{T}(\sigma,z)\ \mathrm{if}\ \sigma\leq t,\\ \ I_{n}\ \ \ \ \ \ \ \ \mathrm{if}\ \sigma>t,\\ \end{matrix}\right.\ \ \ \text{and}\ \ \tilde{v}(\sigma,z)=\left\{\begin{matrix}v(\sigma,y)\ \mathrm{if}\ \sigma\leq t,\\ \ 1\ \ \ \ \ \ \mathrm{if}\ \sigma>t,\\ \end{matrix}\right.$
where $A^{T}$ is the transpose of matrix $A$. Then, $\tilde{v}$ is a nonnegative weak solution to the equation

### v3

@@ 6.1. The Cauchy problem and the generalized fundamental solution
Step 1: regularizing coefficients of $A$: We follow [AN25]. Let $\theta\in\mathcal{D}(\mathbb{R})$ a nonnegative function with $\int_{\mathbb{R}}\theta(t)\mathrm{d}t=1$. For all $p\geq 1$, let $\theta_{p}(t)=p\theta(pt)$ be the associated mollifying sequence. We set $A_{p}(t,x):=(\theta_{p}\star A(\cdot,x))(t)$, i.e., we mollify the matrix-valued function A in the time variable only. For all $p\geq 1$ and $t\in\mathbb{R}$, we set
@@
[…]
@@
Combining Caccioppoli inequality in Lemma 4.4, a weighted Sobolev inequality [HKM06, Theorem 15.26] and the Moser’s iteration principle, we have the following $L^{\infty}$-estimate on nonnegative local weak solutions. For a proof, one can follow the classical scheme or see [ISH99, Proposition 2.1] with lower order coefficients equal to zero.
@@
Moreover, $u\in C([0,\mathfrak{T}];L^{2}_{\omega}(\mathbb{R}^{n}))$ with $u(0)=\psi$, $\lim_{t\to\infty}u(t)=0$ if $\mathfrak{T}=\infty$ (by convention, set $u(\infty)=0$), $t\mapsto\|u(t)\|^{2}_{2,\omega}$ is absolutely continuous on $0,\mathfrak{T}]$ and we can write the energy equalities. Furthermore, we have $(-\Delta_{\omega})^{\alpha/2}u\in L^{r}((0,\mathfrak{T});L^{2}_{\omega}(\mathbb{R}^{n}))$ for any $\alpha\in(0,1]$ with $r=\frac{2}{\alpha}\in[2,\infty)$ with
@@
Setting $g=0$, the case $\mathfrak{T}<\infty$ reproves [AN25, Theorem 1.2], and we even have a larger uniqueness space.
@@ 6.2. Additional properties using regularity theory for weak solutions
Finally, if $\mathcal{O}=\mathbb{R}^{n}$ and $\infty<\mathfrak{T}$, then for all $t>s>\mathfrak{T}$ and $x,y\in\mathbb{R}^{n}$, one has
@@
Using (6.11), we have, for all $0\leq i\leq N-1$, $u(\tau_{i},z_{i})\leq Cu(\tau_{i+1},z_{i+1})$. Thus, by iterating, we get
@@
for some $\delta>0$ depending only on the structural constants, all $t>s$ and almost every $x,y,h\in\mathbb{R}^{n}$ such that $2\left|h\right|\leq(t-s)^{1/2}+\left|x-y\right|$. For proofs and more details, we refer to [AN25] and [TRU68].
@@
[…]
@@
[…]
where $A^{T}$ is the transpose of matrix $A$. By using the $L^{2}_{\omega}(\mathbb{R}^{n})$-valued continuity of $v$, we see that $\tilde{v}$ is a nonnegative weak solution to the equation